OpenStudy (anonymous):

Define the following functions: f(x)=|x| g(x)=|x-3| f+g and f-g and f*g and f/g and g/f

OpenStudy (anonymous):

guys help me plz

hartnn (hartnn):

we know that |x| =x when x>= 0 , and |x|=-x when x<0 similarly, |x-3| = x-3 when x >=3 |x-3| = -(x-3) when x<3 __________________________________________________________ so we consider these intervals, x< 0 where |x| = -x and |x-3| = -(x-3) then f+g = -x-x+3 = -2x+3 in the interval x<0 __________________________________________________________ next interval is x>=0 but x<=3 here, |x| =x , |x-3| = - (x-3) here f+g = x- x+3 = 3 ___________________________________________________________ finally last interval is x>3 when both, |x| = x , and |x-3| = x-3 f+g = x+x-3 = 2x-3 ___________________________________________________________ so, in all, your f+g will be ,